# How Many Van Goghs Does It Take to Van Gogh? Finding the Imitation Threshold

### v1

@@ 7 Analysis: Investigating Outliers
The imitation score plots in the previous section, while showcasing a clear trend, have several outliers. In this section, we analyze such outliers and present examples in Figure 6 (additional outliers can be found in Figures 29 and 30 in the Appendix).
@@
Low Image Counts and High Imitation Scores. Figure 6(a) shows an example of such a case: Thandiwe Newton’s image count is 172 in LAION2B-en, lower than the imitation threshold for celebrities: 364. However, her imitation score of 0.26 is much higher than those of neighboring celebrities with similar image counts (with scores of 0.01 and 0.04). Further investigation reveals that Thandiwe Newton is also known as Thandie Newton. Since this alias may also be used to describe her in captions, MIMETIC2 may have underestimated her image counts. We repeat the process for estimating the image counts with the new alias, and find that Thandie Newton appears in 12,177 images, bringing the cumulative image count to 12,349, which significantly surpasses the established imitation threshold. The two aliases, whose total image count is considerably higher than the imitation threshold, differ by only a single letter and are similarly represented by the model’s encoder (cosine similarity of 0.96), which explains the high imitation score. We find that most of the celebrities from the first kind of outliers are also known by other names which lead to underestimating their image counts. For example, Belle Delphine (394 images) also goes by Mary Belle (310 images, for a total of 704, and DJ Kool Herc (492 images) also goes by Kool Herc (269 images, for a total of 761).
The aliases explanation also largely explains the outliers in art style imitation. For instance, artist Gustav Adolf Mossa (19 images) also goes by just Mossa (15850 images), artist Nicolas Toussaint Charlet (78 images) also goes by just Nicolas Toussaint (533 images), and artist Wilhelm Von Kaulbach (81 images) also goes by Von Kaulbach (978 images). See Figures 31 and 32 in the appendix for the real and generated images of these artists.
High Image Counts and Low Imitation Scores Several celebrities have higher image counts than the imitation threshold, but low imitation scores. Unlike the previous case, we were unable to find a common cause that explains all these outliers. However, we find explanations for specific cases. For example, a staggering proportion of the training images for such celebrities have multiple people in them. For example, out of the 706 total images of Cacee Cobb, only 6 images have her as the only person in the image (see Figure 6(b)). Similarly, out of 1,296 total images of Sofia Hellqvist, only 67 images have her as the only person and out of the 472 total images of Charli D’ Amelio, only 82 images have her as the only person. We hypothesize that having multiple concepts in an image impedes the proper mapping of the concept’s text embedding to its image embedding, which can explain the low imitation score for these concepts. We leave it to future work to further study the connection between the number of concepts in an image and model’s ability to imitate these concepts.
@@ 8 Discussion and Limitations
Equal Effect Assumption. An assumption in the formulation of MIMETIC2 is that every image of a concept contributes equally to the learning of the concept. Although this is a standard assumption in sample complexity works (Valiant, 1984; Udandarao et al., 2024; Wen et al., 2024; Wang et al., 2017), not all images are created equal. While analyzing celebrities’ images for instance, we often find that individuals whose images are mostly close-ups of a single person have a higher imitation score than individuals whose images are cluttered by multiple people, since concept-centered images enhance their learnability. We hope to investigate this assumption in future work.
Factors Affecting the Imitation Threshold. In this work we attribute the imitation of a concept to its image count. However, image count – although a crucial factor – is not the only factor that affects imitation. Several other factors like image resolution, image diversity, alignment between images and their captions, the variance between images of a concept may affect imitation.
Several training time factors like the optimization objective, learning schedule, training data order, model capacity, model architecture also affect the imitation threshold. We discuss the difference in the imitation thresholds of SD1.1, SD1.5 and SD2.1 is attributed to the difference in their text encoders. SD1.1 and SD1.5 use CLIP model Radford et al. (2021) as their text encoder and SD2.1 uses OpenCLIP Ilharco et al. (2021) as its text encoder. Note that while these may impact the behavior of the model, our work is interested in a particular model-data pair, for which we investigate. While we conduct experiments for most models whose training data is available, we do not claim that our results would generalize to other models or datasets, and hope that in future more such models and datasets are open-sourced, so that we can test the generalizability of our imitation thresholds. Note that, MIMETIC2 finds the imitation threshold for the pretraining regime. We hope to investigate the thresholds when a model is finetuned on one or a set of concepts in future work.
Pretraining vs. Finetuning In this work we estimate the imitation threshold of a model that was pretrained on a large dataset. That is, we seek the minimum number of a concept’s images a model has to be trained on to imitate that concept. Note that our question is in contrast to other works that aim to imitate a concept from a minimal number of examples using targeted tuning, such as Dreambooth (Ruiz et al., 2023) and Imagic (Kawar et al., 2023). Our approach and questions are targeted towards copyright and privacy violations of vanilla training of a text-to-image model.
@@ 9 Conclusions
Text-to-image models can imitate their training images (Somepalli et al., 2023a; Carlini et al., 2023a; Somepalli et al., 2023b). This behavior is potentially concerning because these models’ training datasets often include copyrighted and licensed images. Imitating such images would be grounds for violation of copyright and privacy laws. In this work, we seek to find the number of instances of a concept that a text-to-image model needs in order to imitate it – the imitation threshold. We posit this as a new problem, Finding the Imitation Threshold (FIT) and propose an efficient method, MIMETIC2, for finding such thresholds. Our method utilizes pretrained models to estimate this threshold for human faces and art styles imitation using three text-to-image models trained on two different pretraining datasets. We find the imitation threshold of these models to be in the range of 200-600 images depending on the setup. As such, on the domains we evaluate in this work, our results indicate that models cannot imitate concepts that appear less than 200 times in the training data. By estimating the imitation threshold, we provide insights on successful concept imitation based on their training frequencies. Our results have striking implications on both the text-to-image models users and developers. These thresholds can inform text-to-image model developers what concepts are in risk of being imitated, and on the other hand, serve as a basis for copyright and privacy complaints.
@@ 11 Reproducibility Statement
We have published all the code used for this work at https://github.com/vsahil/MIMETIC-2.git and included the README.md file that provides instructions to execute each step of the algorithm.
@@ Acknowledgments
We would like to thank Shauli Ravfogel, Moni Shahar, Amir Feder, Shruti Joshi, and Soumye Singhal for discussion and providing feedback on the draft. In addition, we would like to thank the annotators of our human experiments: Miki Apfelbaum, and the Angle icon by Royyan Wijaya from the Noun Project (CC BY 3.0). A part of this work was supported using Amazon Research Award (ARA) to Chirag Shah.
@@ Appendix A Additional Related Work
Somepalli et al. (2023a); Carlini et al. (2023a) demonstrated that diffusion models can memorize and imitate duplicate images from their training data (they use ‘replication’ to refer to this phenomenon). Casper et al. (2023) corroborated the evidence by showing that these models imitated art styles of 70 artists with high accuracy (as classified by a CLIP model) when prompted to generated images in their styles (a group of artists also sued Stability AI claiming that their widely-used text-to-image models imitated their art style, violating copyright laws Saveri & Butterick (2023)). However, these works did not study how much repetition of a concept’s images would lead the model to imitate them. Studying this relation is important as it serves to guide institutions training these models who want to comply with copyright and privacy laws.
@@
Several works proposed to mitigate the negative impacts of text-to-image models. Shan et al. (2023) proposed GLAZE that adds imperceptible noise to the art works such that diffusion models are unable to imitate artist styles. A similar approach was proposed to hinder learning human faces (Shan et al., 2020). Wang et al. (2024) proposed adding noise to training images, which can be used to detect if a model has been trained on those images. Lu et al. (2024) propose pushing the generated images away from the distribution of training images to minimize mitigation. Kumari et al. (2023); Gandikota et al. (2024) proposed algorithms to remove specific styles, explicit content, and other copyrighted material learned by text-to-image models. On a related note, Xie et al. (2024) proposed Diffusion-ReTrac that finds training images that most influenced a generated image, and thereby provide a fair attribution to training data contributors.
Appendix B Validity of the Assumptions
@@
Table 4: Average difference in the imitation scores for concepts whose image counts differ by less than 10. The difference in the imitation scores are very close to 0, providing the empirical validation the distribution invariance assumption.
@@
In this section, we empirically test the statistical validity of the assumptions we make in Section 3. The first assumption states that there is a distributional invariance across concepts. If this is true, then the imitation scores of two concepts (from the same domain) whose image counts are similar, should also be similar to each other. To test whether this is empirically true for the domains we experiments with, we measure the difference in the imitation scores for concepts whose image counts differ by less than 10 images and report the difference averaged over all such pairs. Table 4 shows that the average difference in the imitation scores for all the pairs, for all the datasets we experiment with, is very close to 0. This provides empirical validation of the distribution invariance assumption.
Appendix C Caption Occurrence Assumption
@@
Table 5: Face count of the ten most popular celebrities in 100K random LAION images. The small percentage of the images we miss shows that our assumption of counting the images where a concept is mentioned in the caption is empirically reasonable.
Celebrity Face Count in 100K images Face Count in Images with Caption Mention Percentage of Missed Images Number of Missed Images
Floyd Mayweather 1 0 0.001% 23K
Oprah Winfrey 2 0 0.002% 46K
Ronald Reagan 6 3 0.003% 69K
Ben Affleck 0 0 0.0% 0
Anne Hathaway 0 0 0.0% 0
Stephen King 0 0 0.0% 0
Johnny Depp 9 1 0.008% 184K
Abraham Lincoln 52 1 0.051% 1.17M
Kate Middleton 34 1 0.033% 759K
Donald Trump 16 0 0.016% 368K
For this purpose, we download 100K random images from LAION2B-en, and run the face detection (used in Section 5) on all images, and count the faces of the ten most popular celebrities in our sampled set of celebrities. Out of the 100K random images, about 57K contain faces. For each celebrity, we compute the similarity between all the faces in the downloaded images and the faces in the reference images of these celebrities. If the similarity is above the threshold of 0.46, we consider that face to belong to the celebrity (this threshold is determined in Appendix E to distinguish if two images are of the same person or not). Table 5 shows the number of faces we found for each celebrity in the 100K random LAION images. We also show the face counts among these images whose captions mention the celebrity. We find that 1) the highest frequency an individual appears in an image without their name mentioned in the caption is 51 (Abraham Lincoln is mentioned once in the caption and he appears a total of 52 times), and 2) the highest percentage of image frequency that we miss is 0.051%, and 3) most of the other miss rates are much smaller (close to 0). Such low miss rates demonstrate that our assumption of counting images when a concept is mentioned in the caption is empirically reasonable.
@@
Appendix D Collection of Reference Images
D.1 Collection of Reference Images for Human Faces Domain
The goal of collecting reference images is to use them to filter the images of the pretraining dataset. These images are treated as the gold standard reference images of a person and images collected from pretraining dataset are compared to these images. If the similarity is higher than a threshold then that image is considered to belong to that person (see Section 5 for details). We describe an automatic manner of collecting the reference images. The high level idea is to collect the images from Google Search and automatically select a subset of those images that are of the same concept (same person’s face or same artist’s art). Since this is a crucial part of the overall algorithm, we manually vet the reference images for all the concepts to ensure that they all contain the same concept.

### v2

@@ 7 Analysis: Investigating Outliers
The imitation score plots in the previous section, while showcasing a clear trend, have several outliers. In this section, we analyze such outliers and present examples in Figure˜6 (more outliers can be found in Figures˜36 and 37 in the Appendix).
@@
Low Image Counts and High Imitation Scores Figure˜6(a) shows an example of such a case: Thandiwe Newton’s image count is 172 in LAION2B-en, lower than the imitation threshold for celebrities: 364. However, her imitation score of 0.26 is much higher than those of neighboring celebrities with similar image counts (with scores of 0.01 and 0.04). Further investigation reveals that Thandiwe Newton is also known as Thandie Newton. Since this alias may also be used to describe her in captions, we underestimate her image counts. We repeat the process for estimating the image counts with the new alias, and find that Thandie Newton appears in 12,177 images, bringing the cumulative image count to 12,349, which significantly surpasses the established imitation threshold. The two aliases, whose total image count is considerably higher than the imitation threshold, differ by only a single letter and are similarly represented by the model’s text encoder (cosine similarity of $0.96$), which explains the high imitation score. We find that most of the celebrities who had a high imitation score while having a low concept frequency are also known by other names which lead to underestimating their image counts. For example, Belle Delphine (394 images) also goes by Mary Belle (310 images), for a total of 704, and DJ Kool Herc (492 images) also goes by Kool Herc (269 images), for a total of 761. The aliases explanation also explains the outliers in art style imitation. For instance, artist Gustav Adolf Mossa (19 images) also goes by just Mossa (15,850 images). See Figures˜38 and 39 in the appendix for the real and generated images in style of these artists.
High Image Counts and Low Imitation Scores Several celebrities have higher image counts than the imitation threshold, but low imitation scores. Unlike the previous case, we were unable to find a common cause that explains all these outliers. However, we find explanations for specific cases. For example, a staggering proportion of the training images for such celebrities have multiple people in them (20% of the outliers of this kind could be explained with this phenomenon). Out of the 706 total images of Cacee Cobb, only 6 images have her as the only person in the image (see Figure˜6(b)). Similarly, out of 1,296 total images of Sofia Hellqvist, only 67 images have her as the only person and out of the 472 total images of Charli D’ Amelio, only 82 images have her as the only person. We hypothesize that having multiple concepts in an image impedes the proper mapping of the concept’s text embedding to its image embedding, which can explain the low imitation score for these concepts. We leave examining this to future work.
@@ 8 Discussion and Limitations
In this work we attribute the imitation of a concept by a model to its frequency in the training data. However, several other factors like image resolution, image diversity, alignment between images and their captions, etc., may affect imitation. Our work is the first to tackle such important question, and as such, we tackle one of the more salient features that contributes to imitation - the concept’s frequency in the training data. In the future, we plan to incorporate other features to improve the accuracy of the imitation threshold. While we hope our work could be used for copyright and privacy claims, our results should be put in content of the assumptions we make (which we find to empirically hold). We further discuss the assumptions we make, and their empirical validity in more detail in Appendix˜B. Note that, MIMETIC2 finds the imitation threshold for the pretraining regime. Different training regimes (such as finetuning on additional data, for multiple epochs) may also affect the imitation threshold, and we hope to investigate such scenarios in future work.
In this work we are interested in finding the estimation threshold of text-to-image models, studying how many images of a given concept a model has to be trained on to be able to generate images of such concepts. We build a system to find such threshold for a given model, dataset, and domain, and show how to estimate such threshold efficiently. We test our method on various domains, datasets, models, and pretraining datasets, and show that the thresholds we estimate are in the same ball park, while exploring some of the variance in our analysis section. Notably, we cannot apriori know how such threshold would generalize to other models, e.g., larger models or models trained on larger datasets that have a different parameter-to-generalization ability. Additional work is needed to study such cases. However, our main contribution in this work is the method (MIMETIC2) that allows to study such threshold and can be applied in future works on additional models.
@@ 9 Conclusions
Text-to-image models can imitate concepts from their training images (somepalli-diffusion-model-copying; carlini2023-extracting-from-diffusion-model; somepalli2023understanding-and-mitigating) This behavior is crucial for learning, but it can also be concerning when such training datasets include copyrighted and private images. Imitating such images would be grounds for violation of copyright and privacy laws. In this work, we seek to find the number of instances of a concept that a text-to-image model has to be trained on in order to imitate it – the imitation threshold. We posit this as a new problem and propose an efficient method, MIMETIC2, for estimating such threshold. Our method uses pretrained models to estimate this threshold for human faces and art styles imitation using four text-to-image models trained on three different pretraining datasets. We find the imitation threshold of these models to be in the range of 200-700 images depending on the setup. By estimating the imitation threshold, we provide insights on successful concept imitation based on their training frequencies. Our results have striking implications on both the text-to-image models users and developers. These thresholds can inform text-to-image model developers what concepts are in risk of being imitated, and on the other hand, serve as a basis for copyright and privacy complaints. Finally, MIMETIC2 also provides insights into the generalization capabilities of text-to-image models.
@@ 11 Reproducibility Statement
We have published all the code used for this work at https://github.com/vsahil/MIMETIC-2.git and included the README.md file that provides instructions to execute each step of the algorithm. We provide our code in the supplementary material. The code includes a detailed README.md file that provides instructions to execute each step of the algorithm. We will open-source all the code upon publication of the paper. We do not provide the original and generated images of individuals and art styles for copyright and privacy concerns.
@@ ACKNOWLEDGMENTS
We would like to thank Shauli Ravfogel, Moni Shahar, Amir Feder, Shruti Joshi, and Soumye Singhal for discussion and providing feedback on the draft. In addition, we would like to thank the annotators of our human experiments: Miki Apfelbaum, and the Angle icon by Royyan Wijaya from the Noun Project (CC BY 3.0). A part of this work was supported using Amazon Research Award (ARA) to Chirag Shah and NSF Grant Numbers IIS-2106937 and IIS-2148367 awarded to Jeff Bilmes.
@@ Appendix A Additional Related Work
somepalli-diffusion-model-copying; carlini2023-extracting-from-diffusion-model demonstrated that diffusion models can memorize and imitate duplicate images from their training data (they use ‘replication’ to refer to this phenomenon). casper2023-DM-artist-imitating-ability corroborated the evidence by showing that these models imitated art styles of 70 artists with high accuracy (as classified by a CLIP model) when prompted to generated images in their styles (a group of artists also sued Stability AI claiming that their widely-used text-to-image models imitated their art style, violating copyright laws artists-sued-SD). However, these works did not study how much repetition of a concept’s images would lead the model to imitate them. Studying this relation is important as it serves to guide institutions training these models who want to comply with copyright and privacy laws.
@@
Several works proposed to mitigate the negative impacts of text-to-image models. Ben-zhao2023-glaze proposed GLAZE that adds imperceptible noise to the art works such that diffusion models are unable to imitate artist styles. A similar approach was proposed to hinder learning human faces (shan2020fawkes-face-privacy). wang2024-detect-illegal-training proposed adding noise to training images, which can be used to detect if a model has been trained on those images. lu2024-dont-reproduce-propulsive propose pushing the generated images away from the distribution of training images to minimize mitigation. kumari2023-erasing-from-DM; gandikota2024-unified-erasing-from-DM proposed algorithms to remove specific styles, explicit content, and other copyrighted material learned by text-to-image models. On a related note, data-attribution-retrak proposed Diffusion-ReTrac that finds training images that most influenced a generated image, and thereby provide a fair attribution to training data contributors.
Appendix B Considerations When Using Our Imitation Thresholds
Our work estimates the imitation thresholds under certain assumptions: distributional invariance between images of different concepts and absence of confounders between the imitation score and image count of a concept. While these assumptions are reasonable (as we discuss in Section˜3 and empirically demonstrate in Section˜C.1) and have been also made in several prior works, they might limit the direct transfer of the imitation thresholds we report to other settings.
Our results also depend on the specific concepts we sampled for each domain, the specific domain embedder we use, and the specific change detection algorithm we use. Changing any of these might alter the imitation thresholds we report. We have performed experiments under several changes to these factors and found our claim to hold. However, we still do not claim our results would generalize to all model/data pairs. Specifically, these are some of the question to consider using our imitation thresholds in a different situation:
  1. 1.
    Is the setting being considered uses the same text-to-image model as we use in our experiments?
  2. 2.
    Is the setting being considered uses the same pretraining dataset as we use in our experiments?
  3. 3.
    Does the setting being considered uses the imitation threshold we report for human face or art style imitation?
  4. 4.
    If our experimental procedure is repeated for a different setting, are the repeated experiments using a similar set of concepts in a domain as we use?
  5. 5.
    If our experimental procedure is repeated for a different setting, is our concept detector or a similar high quality concept detector is being used?
Finally, while we hope our work could be used for copyright and privacy claims, our results should be put in context of the assumptions we make in (Section˜3)
Appendix C Verification of the Assumptions
We empirically test the validity of the three assumptions we make in Section˜3, and find they all hold in practice. To test the invariance assumption we calculate the imitation scores difference of all concepts pairs, whose image counts differ by less than 10 images, for all concepts. We find that the imitation scores difference for such pairs, on average, is 0.003 (averaged across such pairs and averaged across our datasets) (Section˜C.1), attesting to the validity of this assumption. Testing the lack of confounders between imitation score and image count, we finetune a text-to-image model on images of concepts that the model has not seen during pretraining (e.g., politicians that became popular after the training data of the model was curated), and measure the imitation threshold for these concepts Section˜C.2. We argue that if the thresholds found in the finetuning experiment and by MIMETIC2 are close, then the assumption about lack of confounders is valid (assuming there is no other confounder affecting all these concepts). We indeed find that the imitation thresholds found in the finetuning experiment are very close to the thresholds found by MIMETIC2 without finetuning the model ($\pm 6\%$), attesting to the validity of this assumption (Section˜6). Lastly, we test the assumption that each image of a concept contributes equally to the learning of a concept, which is commonplace in the sample complexity literature (pac-learning-theory; wang2017-equal-effect-assumption). We finetune a text-to-image model with a set of randomly sampled images of a concept multiple times. We report the variance in the imitation scores across these finetuning experiments and find the imitation scores variance to be small ($\pm 0.2\%$ of the imitation score), validating this assumption (see more details in Section˜C.3).
To summarize, we empirically verify all the assumptions we make in this work, and find them to hold in practice.
C.1 Validity of the invariance assumption
@@
Table 4: Average difference in the imitation scores for concepts whose image counts differ by less than 10. The difference in the imitation scores are close to 0, empirically validating the distribution invariance assumption.
@@
In this section, we empirically test the statistical validity of the assumptions we make in Section˜3. The first assumption states that there is a distributional invariance across concepts. If this is true, then the imitation scores of two concepts (from the same domain) whose image counts are similar, should also be similar to each other. To test whether this is empirically true for the domains we experiment with, we measure the difference in the imitation scores for concepts whose image counts differ by less than 10 images and report the difference averaged over all such pairs. Table˜4 shows that the average difference in the imitation scores for all the pairs, for all the datasets we experiment with, is very close to 0 (also see Figure˜7 for the distribution of the differences in the score). This provides empirical validation of the distribution invariance assumption.
Figure 7: This figure shows the distribution of differences in imitation scores for concepts whose image counts differ by less than 10 images. We use KDE for this estimation. Since the differences are very small for all the four domains we experiment with, we can argue that our assumption about distributional invariance is valid.
C.2 Validity of the assumption about absence of confounders
To test the absence of confounder, we conduct a finetuning experiment in which we finetuned a pre-trained text-to-image model (SD1.4) on images of concepts the model has not seen during pretraining. We do this by finding politicians who recently got popular, after the SD1.4 model was open-sourced. We finetune SD1.4 on increasing number of images of these politicians, starting with 50 images (much below the imitation threshold found in the non-finetuning setup) and going upto 800 images (above the imitation threshold found in the non-finetuning setup). During the finetuning process, the images of the concepts were mixed with other 10,000 images taken randomly from LAION-2B-en dataset. This was done in order to ensure that our finetuning setup closely resembles the original training setup of SD1.4 where images of a concept would be naturally interspersed with other images in the dataset. We finetune on the full dataset of 10,000 images for 1 epoch with a learning rate of 5e55e-5. We use the following concepts in this experiment: Averie Bishop, Sophie Wilde, Javier Milei, Hashim Safi al-Din, Shyam Rangeela, Akshata Murty, Hana-Rawhiti Maipi-Clarke, Sébastien Lecornu, Vivek Ramaswamy, Raghav Chaddha.
For the SD1.4, the imitation threshold for the non-finetuning setup is found to be at 234 images (Figure˜19) and the average imitation scores for the politicians on the right hand side of the threshold is 0.26. We report the results in Table˜5, which shows that the imitation score remains lower than 0.26 when 50, 100, 150, or 200 images are used to finetune the model. The imitation score reaches this value of 0.26, only when finetuned with 250 or more images. This imitation threshold of 250 in the finetuning experiment is very close to the imitation threshold MIMETIC2 finds for SD1.4 without requiring to finetune the model (234 as shown in Figure˜19).
Note that since in this experiment we are directly intervening on the image count of each concept, there is no effect of any confounder whatsoever even if they were originally present. And the closeness of the imitation threshold in the finetuning and the non-finetuning setup alludes to the absence of confounders, thus validating our assumption. Note that this finetuning experiment is also the optimal experiment for finding the imitation threshold, that we described in Section˜3. Based on these results, we can also claim that imitation thresholds MIMETIC2 finds are close to the ground-truth thresholds that the optimal experiment would find.
Table 5: The imitation scores of the concepts as a text-to-image model, SD1.4, is finetuned on increasing number of images of a concept that it has not seen during pretraining. We report imitation scores averaged over 10 such concepts (politicians). We notice that the imitation scores have values 0.26 or greater when the model is finetuned on 250 or more images of a politician. Critically, this is the average imitation score for politicians on the right hand side of the imitation threshold in Figure˜19 (which is the imitation of politicians for SD1.4), indicating that 250 images or more are required for imitation in this experiment.
# Finetuning Images Imitation Score After Finetuning (Avg. over 10 politicians)
50 0.07
100 0.16
150 0.19
200 0.21
250 0.26
300 0.27
400 0.29
500 0.31
600 0.32
700 0.36
800 0.34
C.3 Validity of the assumption about equal image contribution
Table 6: Mean and std. deviation of the imitation scores when we finetuned a pre-trained SD1.4 model on 400 images of five concepts. We sample the images used for finetuning from a pool of images of these concepts. We repeat this process 20 times, and measure the mean and std. deviation in the imitation scores. The low std. deviation in the imitation scores indicate that most images contribute equally to the learning of the concept, thus validating our assumption.
Concept Mean of Imitation Scores Std. Deviation of Imitation Scores
Javier Milei 0.36 0.02
Hashim Safi al-Din 0.38 0.03
Shyam Rangeela 0.31 0.03
Akshata Murty 0.30 0.03
Hana-Rawhiti Maipi-Clarke 0.30 0.02
We conduct another finetuning experiment to test the validity of our third assumption (which assumes that all images of a concept contribute equally) to its learning. In this experiment, we finetune a pre-trained SD1.4 model on images of concepts that the model has not seen during pre-training (these are the same concepts we used for the finetuning experiment in Section˜C.2). Specifically, in this experiment we sample 400 images of these politicians from a random pool of images 20 times, and finetune a SD1.4 model with each of the sampled set of images. For each of the 20 finetuned models, we then measure the imitation scores for these concepts, and report the mean and standard deviation in the scores (also see Figure˜9 for the distribution of the imitation scores) Table˜6 shows these values, and the low standard deviation in the imitation scores indicate that most images contribute equally to the learning of the concept, thus validating our assumption.
Figure 8: Distribution of imitation scores for the concepts we finetuned on. We use KDE for this estimation.
Figure 9: This figure shows the distribution of imitation scores for concepts we finetuned on using different subsets of training images. We use KDE for this estimation. Since the imitation scores are similar for different subsets of training images for all concepts (demonstrated by narrow curves), we can argue that our assumption of each image contributing equally to model learning is valid.
Appendix D Caption Occurrence Assumption
@@
Table 7: We estimate the percentage of the images that our approach misses due to the assumption that a concept only occurs if it is mentioned in the caption. The small percentage of the images we miss (rightmost column) shows that our assumption of counting the images where a concept is mentioned in the caption is empirically reasonable.
Celebrity Face Count in 100K images Face Count in Images with Caption Mention Missed Images (%)
Floyd Mayweather 1 0 0.001%
Oprah Winfrey 2 0 0.002%
Ronald Reagan 6 3 0.003%
Ben Affleck 0 0 0.0%
Anne Hathaway 0 0 0.0%
Stephen King 0 0 0.0%
Johnny Depp 9 1 0.008%
Abraham Lincoln 52 1 0.051%
Kate Middleton 34 1 0.033%
Donald Trump 16 0 0.016%
For this purpose, we download 100K random images from LAION2B-en, and run the face detection (used in Section˜4) on all images, and count the faces of the ten most popular celebrities in our sampled set of celebrities. Out of the 100K random images, about 57K contain faces. For each celebrity, we compute the similarity between all the faces in the downloaded images and the faces in the reference images of these celebrities. If the similarity is above the threshold of 0.46, we consider that face to belong to the celebrity (this threshold is determined in Appendix˜F to distinguish if two images are of the same person or not). Table˜7 shows the number of faces we found for each celebrity in the 100K random LAION images. We also show the face counts among these images whose captions mention the celebrity. We find that 1) the highest frequency an individual appears in an image without their name mentioned in the caption is 51 (Abraham Lincoln is mentioned once in the caption and he appears a total of 52 times), and 2) the highest percentage of image frequency that we miss is 0.051%, and 3) most of the other miss rates are much smaller (close to 0). Such low miss rates demonstrate that our assumption of counting images when a concept is mentioned in the caption is empirically reasonable.
@@
Appendix E Collection of Reference Images
E.1 Collection of Reference Images for Human Faces Domain
The goal of collecting reference images is to use them to filter the images of the pretraining dataset. These images are treated as the gold standard reference images of a person and images collected from pretraining dataset are compared to these images. If the similarity is higher than a threshold then that image is considered to belong to that person (see Section˜4 for details). We describe an automatic manner of collecting the reference images. The high level idea is to collect the images from Google Search and automatically select a subset of those images that are of the same concept (same person’s face or same artist’s art). Since this is a crucial part of the overall algorithm, we manually vet the reference images for all the concepts to ensure that they all contain the same concept.